# LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Tasks

### v1

@@ 5.3 Experimental Results
LLaMA-2-7B as base model LR. MR. HR. Avg.
@@
LinguaLIFT (MetaMath) 41.2 44.3 45.5 43.2
LinguaLIFT (OpenMath2) 55.4 61.2 62.5 58.9
Table 1: Results on the MMWP benchmark. "LR." denotes the average performance for low-resource languages, "MR." for medium-resource languages, and "HR." for high-resource languages. The bold text highlights the highest score among systems. The dagger symbol (†) indicates results obtained using officially released models, while the double dagger symbol (‡) denotes results from our local implementation. The LLM and extra data used in the third baseline category are indicated in brackets.
LinguaLIFT demonstrates significant improvements across low-resource languages on MMWP.
Experimental results on the MMWP test set, grouped by language resource levels, are presented in Table 1. We elaborate on six key observations from the results: (1) Mono-SFT models exhibit substantial performance drops in low-resource languages and slight drops in medium-resource languages. (2) Multi-SFT models, while outperforming their monolingual counterparts in low-resource settings, still show a considerable performance gap between low-resource and high-resource languages. (3) Models leveraging external translation systems or pre-trained multilingual models achieve limited generalization to unseen languages, and they still lag behind in performance compared to high-resource languages. (4) MindMerger, retrained using additional open-source parallel data in low-resource languages for a fair comparison, performs well in high-resource languages while hardly improving low-resource tasks, demonstrating its limited applicability in low-resource scenarios. (5) LinguaLIFT significantly enhances low-resource reasoning performance, outperforming all competitive baselines. (6) Incorporating advanced English reasoning datasets (e.g., OpenMathInstruct-2) further enhances low-resource language reasoning, highlighting the importance of adapting to evolving, high-quality data for improving low-resource language reasoning.
These results highlight the importance of comprehensively evaluating multilingual reasoning models across diverse languages and further validate LinguaLIFT’s effectiveness in enhancing reasoning performance for low-resource languages.
LinguaLIFT enhances LLM performance in low-resource languages on MGSM and MSVAMP test sets.
@@
Translate-En-MetaMath 39.6 55.1 49.4 47.4 52.3 50.9
LangBridge-MetaMath 38.4 51.9 47.0 43.9 54.5 51.4
MindMerger-Soft-MetaMath 53.1 57.9 56.2 52.7 60.6 58.2
Translate-En-OpenMath2 41.3 63.8 55.6 51.9 62.2 59.1
LangBridge-OpenMath2 42.6 60.8 54.2 47.9 64.4 59.5
MindMerger-Soft-OpenMath2 60.5 67.5 65.0 63.4 74.2 70.9
@@
AlignIFT-MetaMath 55.4 58.8 57.6 56.1 60.6 59.3
AlignIFT-OpenMath2 63.8 66.5 65.5 67.2 74.3 72.2
Table 2: Results on MGSM and MSVAMP datasets. "LR" denotes the average multilingual performance for low-resource languages, while "HR" represents the performance for high-resource languages. The bold text highlights the highest score among systems. The dagger symbol (†) indicates results obtained using officially released models, while the double dagger symbol (‡) denotes results from our local implementation.
Experimental results on the other two widely used multilingual reasoning benchmarks MGSM and MSVAMP are presented in Table 2. These findings are generally consistent with those from the MMWP test sets. LinguaLIFT notably outperforms previous competitive baselines in low-resource languages on both MGSM and MSVAMP test sets, highlighting its strong generalization to various out-of-domain reasoning tasks and robust performance across low-resource language mathematical reasoning tasks.
Versatile Adaptation Across Various Tasks with LinguaLIFT.
@@
Table 3: Results on XNLI and X-CSQA datasets. "LR" denotes the average performance for low-resource languages, and "HR" denotes high-resource languages. The bold text highlights the highest score among systems. The asterisk symbol (*) indicates results are taken from the published results of Huang et al. (2024).
As shown in Table 3, LinguaLIFT outperforms all baselines on language understanding tasks, such as XNLI and X-CSQA, demonstrating its efficiency in adapting various tasks than previous work, as it only needs English-only instruction data while others need to translate instruction data to target language constantly. These results highlight LinguaLIFT’s versatility in generalizing beyond mathematical reasoning, making it a universal and effective approach for enhancing low-resource language task performance.
LinguaLIFT Generalizes well on High-Resource Tasks.
Although LinguaLIFT is designed for low-resource languages, it also boosts performance in high-resource languages. LinguaLIFT outperforms several competitive baselines on mathematical reasoning tasks and multilingual understanding tasks, as shown in Tables 1, 2 and 3. These results highlight LinguaLIFT’s ability to improve both reasoning and comprehension across low- and high-resource languages.
(a) Tow-stage Training
(b) Trainable Modules
Figure 3: Ablation experiments of two-stage training and trainable modules on MGSM. LR., HR., and Avg. represent the average accuracy across low-resource languages, high-resource languages, and all languages, respectively.
Overall, LinguaLIFT demonstrates its effectiveness in enhancing low-resource language capabilities without relying on multilingual instruction data. It established itself as a universally applicable and strong potential method for LLMs to advance low-resource language performance across various tasks. More experimental results are provided in Appendix C.
@@ 6 Analysis
6.1 Two-Stage Training Ablation Studies
Training Stage Ablation
As shown in Figure 3a, removing the Language-Align stage leads to a noticeable performance drop in low-resource languages, demonstrating the essential of the Language-Align stage in enhancing low-resource language tasks. Removing the Task-Transfer stage causes substantial performance degradation across all languages. These results suggest the necessity of the two stages. Further details are provided in Appendix D.1.
@@ 6.1 Two-Stage Training Ablation Studies
Trainable Modules Ablation
As illustrated in Figure 3b, training the LLM during the Language-Align stage reduces performance on high-resource language reasoning. Freezing LLM training in the Task-Transfer stage significantly degrades reasoning performance across all languages. Additionally, training the language alignment layer during the Task-Transfer stage harms performance on low-resource language reasoning. These results indicate the benefit of first training the language alignment layer, followed by LLM fine-tuning. Further details are available in Appendix D.2.
6.2 Analysis of Code-Switched Tuning
Impact of Code-Switched Ratio to Multilingual Alignment and Reasoning Tasks.
@@ 6.2 Analysis of Code-Switched Tuning
We report top-1 retrieval accuracy and reasoning accuracy on Tatoeba (Artetxe and Schwenk, 2019) and MGSM (Shi et al., 2023) to indicate language alignment degree and reasoning performance, respectively, as shown in Figure 4. Multilingual alignment improves as the code-switch ratio increases, especially for low-resource languages, and reasoning performance also increases. Notably, at an 80% code-switch ratio, the model’s reasoning performance reaches comparable performance to those trained with parallel corpora.
Impact of the Part-of-Speech in Code-Switched to Reasoning tasks.
As shown in Figure 6, among the individual part-of-speech (POS) groups, substituting nouns significantly affects the model’s reasoning ability. Syntactic structures involving subject-verb and prepositional phrase combinations outperform other POS combinations, highlighting the importance of core arguments (subjects and verbs) and their relations (prepositions) for capturing key relationships in reasoning tasks. These structures are sufficient for the model to generalize reasoning across languages. In contrast, adjective-adverb and auxiliary-conjunction combinations perform the worst, suggesting that modifiers are less critical for reasoning tasks. More details are provided in Appendix D.3
6.3 LinguaLIFT better aligns low-resource languages
For each low-resource language in the MMWP benchmark, we selected 100 texts with equivalent meaning from the Flores-101 dataset (Goyal et al., 2022). We obtained these input mean pooling representations from different methods and visualized them using T-SNE (van der Maaten and Hinton, 2008). As shown in Figure 5, the LLM embeddings of low-resource languages are distinct from English, indicating the challenges of understanding and transferring knowledge to these languages. In the case of MindMerger, some low-resource language representations closely overlap with English, while others remain isolated, demonstrating its limitations in scenarios involving a wide range of low-resource languages. In contrast, LinguaLIFT aligns low-resource language representations more closely with English, enabling better transfer of reasoning capabilities from English instruction data, thereby resulting in improved reasoning performance.
@@ 6.4 Supplementary Experiments
We conducted several supplementary experiments, including the quantitative analysis of the correlation between multilingual alignment and reasoning performance (Appendix E.1), the analysis of language transferability in language families and writing systems (Appendix E.2), the adaptation of different types and scales of LLMs (Appendix E.3, the selection of the language alignment layer and pre-trained multilingual encoder (Appendix E.4, E.5), the incorporation of multilingual instruction data into LinguaLIFT (Appendix E.6), and zero-shot CoT examples in mathematical reasoning tasks (Appendix E.7).
@@ 7 Conclusion
This paper introduced LinguaLIFT, a novel two-stage instruction tuning framework that enhances low-resource language tasks without relying on parallel corpora or multilingual instruction data. Additionally, we introduced MMWP, a multilingual benchmark spanning 21 low-resource, 17 medium-resource, and 10 high-resource languages, to comprehensively evaluate multilingual mathematical reasoning tasks. Experiments on the MMWP and other widely used benchmarks demonstrate its effectiveness in advancing low-resource language tasks and further alleviating the performance gap between high-resource and low-resource language tasks in LLMs.
@@
While our experimental results demonstrate that the proposed two-stage instruction tuning method significantly improves low-resource language reasoning and understanding tasks, it does require a certain level of computational resources. Specifically, the need for a moderately larger pre-trained multilingual encoder and full fine-tuning of the LLM may impose a computational burden. Additionally, we have not explored parameter-efficient fine-tuning (PEFT) methods, so the effectiveness of our approach in PEFT settings remains untested. Future work could focus on developing more efficient and lightweight tuning strategies for LinguaLIFT to reduce the computational limits.
@@ A.1 Collection of the Alignment Lexicons
To better construct multilingual alignment lexicons for low-resource language tasks, we leveraged the Unsupervised Bilingual Lexicon Induction (UBLI) (Zhang et al., 2017; Lample et al., 2018; Dou et al., 2018; Artetxe et al., 2019; Li et al., 2023b), which has been proven effective in inducing word translation pairs by aligning independently trained word embeddings in two languages.
Initially, we tokenized the text using the Spacy library’s tokenization function [^5] https://spacy.io/. The resultant word set comprised all words, barring specific named entities, numbers, and date tokens. This step is crucial in ensuring our focus on frequent and pertinent terms that are likely to hold significance in task-specific domains.

### v2

@@ 5.3 Experimental Results
LLaMA-2-7B as base model Use LR. MR. HR. Avg.
Multi.
@@
LinguaLIFT (MetaMathQA) 41.2 44.3 45.5 43.2
[…]
LinguaLIFT (OpenMathInstruct-2) 55.4 61.2 62.5 58.9
[…]
Table 1: Results on the MMWP benchmark, where “LR.”, “MR.”, and “HR.” denote mean accuracy for low-, medium-, and high-resource languages, respectively. “Use Multi.” denotes whether multilingual or parallel corpus is utilized. Bold values indicate the highest score across systems. $\dagger$ marks results obtained using officially released models, $\Diamond$ denotes the LoRA (Hu et al., 2022) implementation, and $\ddagger$ signifies results from our local implementation. The LLM and extra data used in the third baseline category are specified in brackets.
LinguaLIFT demonstrates significant improvements across low-resource languages.
Table 1 presents experimental results on the MMWP test set, grouped by language resource levels. We highlight seven key findings: (1) Mono-SFT models exhibit substantial performance drops in low-resource languages. (2) Multi-SFT models outperform monolingual counterparts in low-resource settings but retain a notable performance gap between low- and high-resource languages. (3) Models using external translation systems or pre-trained multilingual models generalize poorly to unseen languages and underperform in low-resource contexts. (4) Even with supplementary low-resource data, MindMerger fails to improve low-resource language reasoning performance. (5) LinguaLIFT substantially improves low-resource reasoning accuracy, surpassing all baselines without requiring multilingual or parallel data. (6) Integrating advanced English reasoning datasets (e.g., OpenMathInstruct-2) further enhances reasoning performance, underscoring the need to adapt to evolving high-quality English reasoning data.
Robust generalization and task versatility.
@@
Translate-En (MetaMath) 39.6 55.1 49.4 47.4 52.3 50.9
LangBridge (MetaMath) 38.4 51.9 47.0 43.9 54.5 51.4
MindMerger-Soft (MetaMath) 53.1 57.9 56.2 52.7 60.6 58.2
Translate-En (OpenMath2) 41.3 63.8 55.6 51.9 62.2 59.1
LangBridge (OpenMath2) 42.6 60.8 54.2 47.9 64.4 59.5
MindMerger-Soft (OpenMath2) 60.5 67.5 65.0 63.4 74.2 70.9
@@
LinguaLIFT (MetaMathQA) 55.4 58.8 57.6 56.1 60.6 59.3
[…]
LinguaLIFT (OpenMathInstruct-2) 63.8 66.5 65.5 67.2 74.3 72.2
[…]
Table 2: Results on MGSM and MSVAMP datasets. The usage of abbreviation, symbol and boldface are the same in Table 1.
Experimental results across four multilingual reasoning benchmarks (Tables 2 and 3) demonstrate LinguaLIFT’s robust generalization capabilities in out-of-domain scenarios. The framework achieves significant improvements over existing baselines in mathematical reasoning tasks (MGSM, MSVAMP) for low-resource languages while exhibiting strong adaptability to cross-domain commonsense reasoning challenges (XNLI, X-CSQA). By excelling in numerical computation and understanding semantics, LinguaLIFT demonstrates its task-agnostic generalization ability across various tasks, effectively transferring learned patterns between different reasoning paradigms and thus enhancing performance in low-resource language reasoning. Furthermore, the model enhances reasoning accuracy even in high-resource languages, underscoring its dual capacity to generalize and adapt across tasks and resource levels.
@@
Table 3: Results on XNLI and X-CSQA datasets. The usage of abbreviation is the same in Table 2. The asterisk symbol (*) indicates results are taken from the published results of Huang et al. (2024).
Overall, LinguaLIFT established itself as a universally applicable and strong potential method for LLMs to advance low-resource language performance across various reasoning tasks. More experimental results are provided in Appendix C.
@@ 6 Analysis
6.1 Two-stage Training Ablation
Figure 3: Ablation study of two-stage training on MGSM, showing average accuracy for low-resource (LR.), high-resource (HR.), and all languages (Avg.).
As shown in Figure 3, ablating the Language Align Stage results in a significant performance decline for low-resource languages, empirically validating its role in cross-lingual knowledge transfer. Similarly, removing the Task Transfer Stage degrades performance across all languages. These results underscore both stages’ mutually dependent contributions to enhancing reasoning tasks. Further details are provided in Appendix D.1.
6.2 Trainable Modules Ablation
Figure 4: Trainable modules ablation on MGSM. The usage of abbreviation is the same in Figure 3.
As shown in Figure 4, training the LLM during the Language Align stage impairs high-resource language reasoning, likely due to catastrophic forgetting of acquired knowledge. Freezing the LLM during the Task Transfer stage similarly degrades performance across all languages. Notably, training the language alignment layer in this stage undermines low-resource reasoning, as it disrupts the pre-established multilingual alignment. These findings necessitate sequential training: first optimizing the alignment layer, followed by LLM fine-tuning. Further details are available in Appendix D.2.
6.3 Multilingual Encoder Ablation
Figure 5: Accuracy (%) of LinguaLIFT across encoders’ scales and types on MGSM’s low-resource reasoning.
Figure 5 evaluates low-resource language reasoning on the MGSM test set using the mT5 series (Xue et al., 2021) across various scales and four pre-trained multilingual encoders: mBERT (Devlin et al., 2019), XLM (Lample and Conneau, 2019), XLM-R (Conneau et al., 2020), and LaBSE (Feng et al., 2022). Performance improved significantly as mT5 scaled from 270M to 2.2B parameters but stagnated beyond this threshold, indicating a plateau at larger scales. Notably, LaBSE, leveraging parallel corpora for robust cross-lingual alignment, outperformed other encoders. More detailed information is provided in Appendix D.3.
6.4 Generalizes Well to Various LLMs
Mistral-7B Llama-2-13B
Methods LR. HR. Avg. LR. HR. Avg.
baseline
Mono-SFT(English-only Task Data)
MetaMath* 30.0 66.5 55.6 8.53 59.7 44.4
Multi-SFT(Multilingual Task Data with Query Translation)
MathOctopus-Parallel* 48.1 51.1 50.2 41.6 47.7 45.8
QAlign-MetaMathQA* 50.7 59.3 56.7 44.7 62.4 57.1
Freezing LLM with External Tools or Models
Translate-En* 53.7 60.1 58.2 44.4 55.5 52.2
LangBridge* 52.4 65.8 61.8 41.3 52.9 49.5
MindMerger-Soft* 56.8 69.1 65.4 57.1 65.1 62.7
our methods
LinguaLIFT (MetaMathQA) 57.7 68.8 65.5 58.7 64.7 62.9
Table 4: Tesults on MGSM across various LLMs. Abbreviation and symbol are the same in Table 1.
We extend the evaluations in Section 5 to diverse LLMs and scales, including Llama-2-13B Touvron et al. (2023) and Mistral-7B (Jiang et al., 2023). Table 4 demonstrates LinguaLIFT’s broader applicability across LLM scales and architectures, with full results in Appendix D.4.
6.5 Analysis of Language Transferability
Figure 6: Reasoning accuracy on MMWP group by language families and writing systems.
As illustrated in Figure 6, LinguaLIFT outperformed all competitive baselines, achieving state-of-the-art transferability across language families and writing systems, particularly in languages sharing English’s syntax or script (e.g., Germanic, Latin). These findings highlight LinguaLIFT as a robust cross-lingual transfer framework, leveraging shared linguistic features to enhance low-resource language reasoning tasks. More details are provided in Appendix D.5.
6.6 Analysis of Code-Switched Ratio
@@
We assess language alignment and mathematical reasoning via top-1 retrieval accuracy on Tatoeba (Artetxe and Schwenk, 2019) and MGSM accuracy, respectively (Figure 7). Higher code-switching ratios enhance multilingual alignment, especially for low-resource languages, with proportional gains in reasoning accuracy. Notably, at an 80% ratio, the model matches the performance of parallel corpus-trained systems, indicating that word-level code-switching suffices for robust alignment. This demonstrates that explicit parallel corpora are unnecessary, as code-switching inherently promotes multilingual understanding and reasoning in low-resource scenarios.
6.7 Analysis of Code-Switched Part-of-Speech
@@
As illustrated in Figure 8, noun substitutions significantly impair the model’s reasoning capabilities among part-of-speech combinations. Syntactic structures involving subject-verb pairs and prepositional phrases demonstrate superior performance compared to other combinations, underscoring the importance of core arguments (subjects, verbs) and their syntactic relations (prepositions) in capturing essential relational information for cross-linguistic reasoning generalization. Conversely, adjective-adverb and auxiliary-conjunction combinations exhibit the lowest performance, indicating that modifiers play a less critical role in reasoning. Further details are available in Appendix D.6.
6.8 Incorporating Multilingual Data
Figure 10: Accuracy (%) on MGSM when LinguaLIFT leverages multilingual instruction and parallel data.
Beyond using code-switched translation data, LinguaLIFT can adapt to the training setting of previous works by integrating real multilingual data, further advancing reasoning performance. As shown in Figure 10, Multi-LinguaLIFT achieves a higher ceiling for low-resource language reasoning, with full experimental details in Appendix D.7.
6.9 Alignment Representation Visualization
For each low-resource language in the MMWP benchmark, we extracted 100 semantically equivalent texts from the Flores-101 dataset (Goyal et al., 2022). We derived mean pooling representations using multiple methods and visualized them via T-SNE (van der Maaten and Hinton, 2008). As shown in Appendix E.1, cross-lingual alignment positively correlates with reasoning transfer: representations closer to English exhibit more substantial reasoning capability generalization. Figure 9 reveals that the LLaMA2-7B’s low-resource language representations are scattered and divergent from English, indicating the challenges in knowledge transfer. MindMerger achieves partial alignment, with some low-resource language representations overlapping English representations; however, others remain isolated, reflecting its instability in diverse low-resource scenarios. In contrast, LinguaLIFT tightly aligns low-resource languages with English, enabling effective reasoning transfer from English instruction data and yielding superior performance.
@@ 6.10 Supplementary Experiments
More supplementary experiments comprising: (1) quantitative correlations between multilingual alignment and reasoning performance (Appendix E.1), (2) language alignment layer selection (Appendix E.2), and (3) zero-shot CoT examples in mathematical reasoning (Appendix E.3).
@@ 7 Conclusion
This paper proposes LinguaLIFT, a novel two-stage instruction tuning framework that enhances low-resource language reasoning without relying on parallel corpora or multilingual instruction data. Additionally, we introduce MMWP, a multilingual benchmark spanning 21 low-resource, 17 medium-resource, and 10 high-resource languages, to comprehensively evaluate multilingual mathematical reasoning tasks. Experiments on the MMWP and other widely used benchmarks demonstrate its effectiveness in advancing low-resource language reasoning and further alleviating the performance gap between high-resource and low-resource language reasoning in LLMs.
@@
While our experimental results demonstrate that leveraging code-switched tuning and cross-lingual transfer can help improve low-resource language reasoning, its implementation entails computational considerations requiring a moderately-size multilingual encoder for cross-lingual alignment, which introduces notable resource demands. Although we employ LoRA for parameter efficiency, this adaptation may underperform full-parameter tuning in specific scenarios. Future work should prioritize developing lightweight adaptation strategies that preserve cross-lingual transfer capabilities while optimizing computational efficiency.
@@ A.1 Collection of the Alignment Lexicons
To better construct multilingual alignment lexicons for low-resource language reasoning, we leveraged the Unsupervised Bilingual Lexicon Induction (UBLI) (Zhang et al., 2017; Lample et al., 2018; Dou et al., 2018; Artetxe et al., 2019; Li et al., 2023b), which has been proven effective in inducing word translation pairs by aligning independently trained word embeddings in two languages.
Initially, we tokenized the text using the Spacy library’s tokenization function [^2] https://spacy.io/. The resultant word set comprised all words, barring specific named entities, numbers, and date tokens. This step is crucial in ensuring our focus on frequent and pertinent terms that are likely to hold significance in reasoning domains.
@@
The accuracy of the word mappings in our code-switch approach is a critical consideration, especially in the absence of parallel data. To demonstrate the feasibility and effectiveness of our unsupervised method, we address this important question from multiple perspectives:
$\bullet$ Methodology Validation: The unsupervised word translation method we adopted is based on MUSE (Lample et al., 2018), which has been thoroughly validated in previous research. As demonstrated in MUSE (Section 4.2), this approach achieves remarkable performance without any parallel data - up to 80% accuracy for high-resource languages and around 50% for low-resource languages when translating to English, which is competitive with supervised methods. Constructing a lexicon in such settings is inherently challenging, yet our method achieves remarkable performance, demonstrating the novelty and effectiveness of our approach.
$\bullet$ Further evaluation using GPT-4o: To evaluate the accuracy of the lexicons constructed in Stage 1, we employed GPT-4o to evaluate the quality of the generated word mappings. The evaluation, using P@5 accuracy, yielded impressive results: (i) High-resource languages: 87.4% accuracy (P@5) and (ii) Low-resource languages: 61.9% accuracy (P@5).
These results further validate the quality of the unsupervised word mappings produced by our method, underscoring the strength of our approach even in low-resource settings.
@@ A.2 Dataset Statistics
In this study, we utilize a variety of datasets for training and evaluation purposes, as detailed in Table 6. The datasets are divided into two main categories: Training Datasets and Evaluation Datasets.
The Training Datasets include data from different sources and cover a range of languages and sample sizes. These datasets are primarily used for training the models:
  • MetaMathQA (Yu et al., 2024): A dataset with 395,000 samples in a single language, primarily used for mathematical question-answering tasks.
  • OpenMathInstruct-2 (Toshniwal et al., 2024): This dataset contains 1,000,000 samples and is also in a single language, focusing on instructional tasks related to mathematics.
  • MGSM8KInstruct (Chen et al., 2023): A multilingual dataset involving 10 languages and 73,599 samples, designed for general-purpose instruction-based tasks.
  • Multi-NLI (Williams et al., 2018): With 392,702 samples, this dataset is in a single language and focuses on natural language inference tasks.
  • QASC (Khot et al., 2020): A dataset with 8,134 samples used for question-answering tasks based on commonsense reasoning in a single language.
  • ARC (Clark et al., 2018): The ARC dataset contains 3,370 samples for answering elementary science questions in a single language.
  • OpenBookQA (Mihaylov et al., 2018): This dataset consists of 4,957 samples specifically designed for open-book question answering in a single language.
  • X-CSQA(En) (Lin et al., 2021): A dataset of 8,888 samples for cross-lingual commonsense question answering in English.
Evaluation Datasets
The Evaluation Datasets are used to assess the performance of the trained models across different tasks and languages:
  • MMWP: A multilingual dataset with 38,400 samples covering 48 languages. It is used for evaluating mathematical word problems.
  • MGSM (Shi et al., 2023): With 11 languages and 2,750 samples, this dataset is used for evaluating general language model performance across multiple languages.
  • MSVAMP (Chen et al., 2023): A dataset for evaluating model performance on visual question-answering tasks, containing 10,000 samples across 10 languages.
  • XNLI (Conneau et al., 2018): This evaluation dataset covers 15 languages and contains 75,150 samples used for cross-lingual natural language inference tasks.
  • X-CSQA (Lin et al., 2021): A dataset with 16,000 samples, spanning 16 languages, designed for cross-lingual commonsense question answering.
The number of languages (#Lang) and number of samples (#Samples) provided in each dataset enable a comprehensive evaluation of models across various linguistic and task-specific domains.